# Stability-Certified On-Policy Data-Driven LQR via Recursive Learning and Policy Gradient

### v1

@@ 4.1 Closed-Loop Dynamics in Error Coordinates
As a preliminary step, system (3) is expressed into suitably defined error coordinates. First, we consider vectorized versions of the matrix updates in (19g)-(19l). To this end, let the new coordinates $H^{\text{vc}}\in\mathbb{R}^{(n+m)^{2}}$ and $S^{\text{vc}}\in\mathbb{R}^{(n+m)n}$ be defined as
$\begin{cases}H_{t}\\ S_{t}\\ \end{cases}\longmapsto\begin{cases}H^{\text{vc}}_{t}:=\mathtt{vec}\left(H_{t}\right)\\ S^{\text{vc}}_{t}:=\mathtt{vec}\left(S_{t}\right).\end{cases}$ (21)
Therefore, (19g)-(19l) can be recast as
$H^{\text{vc}}_{t+1}\!$ $=\!\lambda H^{\text{vc}}_{t}\!+\!\mathtt{vec}\left(\!\begin{bmatrix}\mathrm{\mathrm{x}}_{t}\\ K_{t}\mathrm{\mathrm{x}}_{t}\!+\!E\mathrm{w}_{t}\end{bmatrix}\!\!\!\begin{bmatrix}\mathrm{\mathrm{x}}_{t}\\ K_{t}\mathrm{\mathrm{x}}_{t}\!+\!E\mathrm{w}_{t}\end{bmatrix}^{\top}\!\right)$ (22e)
$S^{\text{vc}}_{t+1}\!$ $=\!\lambda S^{\text{vc}}_{t}\!+\!\mathtt{vec}\left(\!\begin{bmatrix}\mathrm{\mathrm{x}}_{t}\\ K_{t}\mathrm{\mathrm{x}}_{t}\!+\!E\mathrm{w}_{t}\end{bmatrix}\!\!\!\begin{bmatrix}\mathrm{\mathrm{x}}_{t}\\ K_{t}\mathrm{\mathrm{x}}_{t}\!+\!E\mathrm{w}_{t}\end{bmatrix}^{\top}\!\!\!\theta^{\star}\!\right)\!\!.$ (22j)
Next, we will inspect (4.1) together with (19b) to provide the steady-state locus (see, e.g., [50, Ch. 12] for a formal definition) when the system is fed with the signal $\mathrm{w}_{t}$, which evolves according to (19a). To this end, set $n_{\chi}:=n+(n+m)^{2}+(n+m)n$ and let $\chi\in\mathbb{R}^{n_{\chi}}$ be defined as
$\chi:=\begin{bmatrix}\mathrm{x}^{\top}&(H^{\text{vc}})^{\top}&(S^{\text{vc}})^{\top}\end{bmatrix}^{\top}.$
Then, using (4.1), the dynamics in (19a)-(19l) can be compactly expressed in the new coordinates as
@@
$\mathcal{A}_{K}(K)$ $:=\begin{bmatrix}{A}_{\star}+{B}_{\star}K&0&0\\ 0&\lambda I&0\\ 0&0&\lambda I\end{bmatrix}$ (24d)
$\phi(\chi,K,\mathrm{w})$ $:=\begin{bmatrix}{B}_{\star}E\mathrm{w}\\ \mathtt{vec}\left(\begin{bmatrix}\mathrm{x}\\ K\mathrm{x}\!+\!E\mathrm{w}\end{bmatrix}\begin{bmatrix}\mathrm{x}\\ K\mathrm{x}\!+\!E\mathrm{w}\end{bmatrix}^{\top}\right)\\ \mathtt{vec}\left(\begin{bmatrix}\mathrm{x}\\ K\mathrm{x}\!+\!E\mathrm{w}\end{bmatrix}\begin{bmatrix}\mathrm{x}\\ K\mathrm{x}\!+\!E\mathrm{w}\end{bmatrix}^{\top}\theta^{\star}\right)\end{bmatrix}.$ (24p)
Notice that to keep the notation light, we use a hybrid notation with $\chi$ on the left-hand side and its (unvectorized) components $(\mathrm{x},H,S)$ on the right-hand side.
System (4.1) together with the exosystem (19a) is a cascade whose steady-state locus can be characterized by the nonlinear map $\chi^{{\textnormal{ss}}}:\mathbb{R}^{n_{\mathrm{w}}}\to\mathbb{R}^{n_{\chi}}$ defined as
$\chi^{{\textnormal{ss}}}(\mathrm{w}):=\begin{bmatrix}\Pi_{\mathrm{x}}\mathrm{w}\\ \Pi_{H}\mathtt{vec}\left(\mathrm{w}\mathrm{w}^{\top}\right)\\ \Pi_{S}\mathtt{vec}\left(\mathrm{w}\mathrm{w}^{\top}\right)\end{bmatrix},$ (28)
where $\Pi_{\mathrm{x}}$, $\Pi_{H}$, and $\Pi_{S}$ are the same as in Theorem 3.2 (see (B.2) and (B) in Appendix B for their explicit definition). Formally, the following lemma holds true.
@@
Let the assumptions of Theorem 3.2 hold true. Consider the map $\chi^{{\textnormal{ss}}}$ defined in (28), the feedback gain $K^{\star}$ solving (2.1), the matrix $F$ as in (3), and the functions $\mathcal{A}_{K}$ and $\phi$ defined in (4.1). Then, it holds
@@
normal-$\blacksquare$
The proof of Lemma 4.1 is provided in Appendix B.
Lemma 4.1 ensures that $\operatorname{col}(\chi^{{\textnormal{ss}}}(\mathrm{w}),\theta^{\star},K^{\star})$ is the steady-state locus of the overall closed-loop system (3). In this regard, we also included condition (30) since it allows us to show that $\theta^{\star}$ is an equilibrium of update (19m) restricted to the case in which $H_{t}$ and $S_{t}$ lie in the steady-state locus. Indeed, when $\chi_{t}=\chi^{{\textnormal{ss}}}(\mathrm{w}_{t})$, the update (19m) reduces to
@@
=$\stackrel{{\scriptstyle(a)}}{{=}}\theta_{t}-\gamma\mathrm{v}_{H}(\Pi_{H},\mathrm{w}_{t})\left(\theta_{t}-\theta^{\star}\right),$
where in $(a)$ we used a property of the vectorization operator [^1] Given any two matrices $X_{1}\in\mathbb{R}^{n_{1}\times n_{2}}$ and $X_{2}\in\mathbb{R}^{n_{2}\times n_{3}}$, it holds $\mathtt{vec}\left(X_{1}X_{2}\right)=(X_{2}^{\top}\otimes I_{n_{1}})\mathtt{vec}\left(X_{1}\right)$.. As for the equilibrium of (19n) when the other states lie on the steady-state locus, it turns out to be $K^{\star}$ since $G(K^{\star},\theta^{\star})=0$.
@@
$\mathrm{z}:=\begin{bmatrix}\theta^{\top}&K^{\top}\end{bmatrix}^{\top}.$
@@
$\begin{cases}\mathrm{w}\\ \chi\\ \mathrm{z}\end{cases}\longmapsto\begin{cases}\mathrm{w}\\ \tilde{\chi}:=\begin{bmatrix}I&0&0\\ 0&\gamma I&0\\ 0&0&\gamma I\end{bmatrix}\left(\chi-\chi^{{\textnormal{ss}}}(\mathrm{w})\right)\\ \tilde{\mathrm{z}}:=\mathrm{z}-\begin{bmatrix}\theta^{\star}\\ K^{\star}\end{bmatrix}.\end{cases}$ (31)
For notational convenience, we will sometimes refer to the components of $\tilde{\mathrm{z}}$ as $\operatorname{col}(\tilde{\theta},\tilde{K})$. Finally, the closed-loop dynamics (3) in the new coordinates (31) reads as
@@
where $\mathcal{A}(\tilde{\mathrm{z}}):=\mathcal{A}_{K}(\tilde{K}+K^{\star})$ and we introduced $h:\mathbb{R}^{n_{\mathrm{z}}}\times\mathbb{N}\to\mathbb{R}^{n_{\mathrm{z}}}$, $g:\mathbb{R}^{n_{\chi}}\times\mathbb{R}^{n_{\mathrm{z}}}\times\mathbb{N}\to\mathbb{R}^{n_{\chi}}$, and $f:\mathbb{R}^{n_{\chi}}\times\mathbb{R}^{n_{\mathrm{z}}}\times\mathbb{N}\to\mathbb{R}^{n_{\mathrm{z}}}$ defined respectively as
$h(\tilde{\mathrm{z}},t)$ $:=\begin{bmatrix}{B}_{\star}\tilde{K}\Pi_{\mathrm{x}}\mathrm{w}_{t}\\ 0\\ 0\end{bmatrix}$ (33d)
$g(\tilde{\chi},\tilde{\mathrm{z}},t)$ $:=\begin{bmatrix}0\\ \gamma\phi(\tilde{\mathrm{x}}+\Pi_{\mathrm{x}}\mathrm{w}_{t},\tilde{K}+K^{\star},\mathrm{w}_{t})\end{bmatrix}$ (33g)
$f(\tilde{\chi},\tilde{\mathrm{z}},t)$ $:=-\begin{bmatrix}(\tilde{H}+H^{{\textnormal{ss}}}_{t})^{\dagger}\left((\tilde{H}+H^{{\textnormal{ss}}}_{t})\tilde{\theta}+(\tilde{H}-\tilde{S})\theta^{\star}\right)\\ G(\tilde{K}+K^{\star},\tilde{\theta}+\theta^{\star})\end{bmatrix}.$ (33j)
Notice that for the sake of readability, in (4.1) we used the shorthands $\tilde{\chi}:=\operatorname{col}(\tilde{\mathrm{x}},\tilde{H},\tilde{S})$ and $\tilde{\mathrm{z}}=\operatorname{col}(\tilde{\theta},\tilde{K})$, we defined $H^{{\textnormal{ss}}}_{t}\in\mathbb{R}^{(n+m)\times(n+m)}$ as
@@
which represents the steady-state value of the state $H_{t}$ and we introduced the error coordinates $\tilde{H}\in\mathbb{R}^{(n+m)\times(n+m)}$ and $\tilde{S}\in\mathbb{R}^{(n+m)\times n}$, as
$\tilde{H}$ $:=\mathtt{unvec}\left(H^{\text{vc}}-H^{{\textnormal{ss}}}_{t}\right)$ (35a)
$\tilde{S}$ $:=\mathtt{unvec}\left(S^{\text{vc}}-\Pi_{S}\mathtt{vec}\left(\mathrm{w}_{t}\mathrm{w}_{t}^{\top}\right)\right).$ (35b)
Some remarks are in order. We point out that with this transformation we obtained a dynamical system with two-time scales as the one described in Appendix A (cf. system (A)). As customary in the context of singularly perturbed systems, we distinguish between (i) the fast dynamics (32a) with state $\tilde{\chi}$, and (ii) the slow one (32b) with state $\tilde{\mathrm{z}}$. Figure 3 shows the mentioned interconnected structure of system (4.1).
Figure 3: Block diagram describing system (4.1).
It is also worth noting that, in this reformulation, the effect of the exogenous/dithering signal $\mathrm{w}_{t}$ has been embedded in the time dependency of $h$, $g$, and $f$. As for the equilibrium manifold, we observe that it holds
@@
for all $t\in\mathbb{N}$. Finally, the matrix $\mathcal{A}(\tilde{\mathrm{z}})$ is Schur for all $\tilde{K}\in\mathbb{R}^{m\times n}$ such that $(\tilde{K}+K^{\star})\in\mathcal{K}$.

### v2

@@ 4.1 Closed-Loop Dynamics in Error Coordinates
As a preliminary step, system (19) is expressed into suitably defined error coordinates. First, we consider vectorized versions of the matrix updates in (19c)-(19d). To this end, let the new coordinates $H^{\text{vc}}\in\mathbb{R}^{(n+m)^{2}}$ and $S^{\text{vc}}\in\mathbb{R}^{(n+m)n}$ be defined as
$\begin{bmatrix}H\\ S\end{bmatrix}\longmapsto\begin{bmatrix}H^{\text{vc}}\\ S^{\text{vc}}\end{bmatrix}:=\begin{bmatrix}\mathtt{vec}\left(H\right)\\ \mathtt{vec}\left(S\right)\end{bmatrix}.$ (22)
Therefore, (19c)-(19d) can be recast as
$H^{\text{vc}}_{t+1}\!$ $=\!\lambda H^{\text{vc}}_{t}\!+\!\mathtt{vec}\left(\!\begin{bmatrix}\mathrm{\mathrm{x}}_{t}\\ K_{t}\mathrm{\mathrm{x}}_{t}\!+\!E\mathrm{w}_{t}\end{bmatrix}\!\!\!\begin{bmatrix}\mathrm{\mathrm{x}}_{t}\\ K_{t}\mathrm{\mathrm{x}}_{t}\!+\!E\mathrm{w}_{t}\end{bmatrix}^{\top}\!\right)$ (23a)
$S^{\text{vc}}_{t+1}\!$ $=\!\lambda S^{\text{vc}}_{t}\!+\!\mathtt{vec}\left(\!\begin{bmatrix}\mathrm{\mathrm{x}}_{t}\\ K_{t}\mathrm{\mathrm{x}}_{t}\!+\!E\mathrm{w}_{t}\end{bmatrix}\!\!\!\begin{bmatrix}\mathrm{\mathrm{x}}_{t}\\ K_{t}\mathrm{\mathrm{x}}_{t}\!+\!E\mathrm{w}_{t}\end{bmatrix}^{\top}\!\!\!\theta^{\star}\!\right)\!\!.$ (23b)
Next, we will inspect (23) together with (19b) to provide the steady-state locus (see, e.g., [53, Ch. 12] for a formal definition) when the system is fed with the signal $\mathrm{w}_{t}$, which evolves according to (19a). To this end, set $n_{\chi}:=n+(n+m)^{2}+(n+m)n$ and let $\chi\in\mathbb{R}^{n_{\chi}}$ be defined as
$\chi:={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\operatorname{col}(\mathrm{x},H^{\text{vc}},S^{\text{vc}}).}$
Then, using (23), the dynamics in (19a)-(19d) can be compactly expressed in the new coordinates as
@@
$\mathcal{A}_{K}(K)$ $:=\begin{bmatrix}{A}_{\star}+{B}_{\star}K&0&\\ 0&{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\lambda I_{({n}+{m})(2{n}+{m})}}\end{bmatrix}$ (25a)
$\phi(\chi,K,\mathrm{w})$ $:=\begin{bmatrix}{B}_{\star}E\mathrm{w}\\ \mathtt{vec}\left(\begin{bmatrix}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\chi_{1}}\\ K{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\chi_{1}}\!+\!E\mathrm{w}\end{bmatrix}\begin{bmatrix}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\chi_{1}}\\ K{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\chi_{1}}\!+\!E\mathrm{w}\end{bmatrix}^{\top}\right)\\ \mathtt{vec}\left(\begin{bmatrix}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\chi_{1}}\\ K{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\chi_{1}}\!+\!E\mathrm{w}\end{bmatrix}\begin{bmatrix}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\chi_{1}}\\ K{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\chi_{1}}\!+\!E\mathrm{w}\end{bmatrix}^{\top}\theta^{\star}\right)\end{bmatrix},$ (25b)
in which $\chi_{1}\in\mathbb{R}^{{n}}$ denotes the first ${n}$ components of $\chi$.
System (24) together with the exosystem (19a) is a cascade whose steady-state locus can be characterized by the nonlinear map $\chi^{{\textnormal{ss}}}:\mathbb{R}^{n_{\mathrm{w}}}\to\mathbb{R}^{n_{\chi}}$ defined as
$\chi^{{\textnormal{ss}}}(\mathrm{w}):=\begin{bmatrix}\Pi_{\mathrm{x}}\mathrm{w}\\ {\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\begin{bmatrix}\Pi_{H}&0\\ 0&\Pi_{S}\end{bmatrix}\mathtt{vec}\left(\mathrm{w}\mathrm{w}^{\top}\right)}\end{bmatrix},$ (26)
where $\Pi_{\mathrm{x}}$, $\Pi_{H}$, and $\Pi_{S}$ are those referred in Theorem 3.3 (see (A.2) and (A.6) in Appendix A for their explicit definition). Formally, the following lemma holds true.
@@
Let the assumptions of Theorem 3.3 hold true. Consider the map $\chi^{{\textnormal{ss}}}$ defined in (26), the feedback gain $K^{\star}$ solving (5), the matrix $F$ as in (17), and the functions $\mathcal{A}_{K}$ and $\phi$ defined in (25). Then, it holds
@@
$\blacksquare$
The proof of Lemma 4.1 is provided in Appendix A.
Lemma 4.1 ensures that $\operatorname{col}(\chi^{{\textnormal{ss}}}(\mathrm{w}),\theta^{\star},K^{\star})$ is the steady-state locus of the overall closed-loop system (19). In this regard, we also included condition (28) since it allows us to show that $\theta^{\star}$ is an equilibrium of update (19e) restricted to the case in which $H_{t}$ and $S_{t}$ lie in the steady-state locus. Indeed, when $\chi_{t}=\chi^{{\textnormal{ss}}}(\mathrm{w}_{t})$, the update (19e) reduces to
@@
$=\theta_{t}-\gamma\mathrm{v}_{H}(\Pi_{H},\mathrm{w}_{t})\left(\theta_{t}-\theta^{\star}\right),$
where we used a property of the vectorization operator [^1] Given any two matrices $X_{1}\in\mathbb{R}^{n_{1}\times n_{2}}$ and $X_{2}\in\mathbb{R}^{n_{2}\times n_{3}}$, it holds $\mathtt{vec}\left(X_{1}X_{2}\right)=(X_{2}^{\top}\otimes I_{n_{1}})\mathtt{vec}\left(X_{1}\right)$. and the definitions of $\mathrm{v}_{H}$ and $\mathrm{v}_{S}$ given in (20) As for the equilibrium of (19f) when the other states lie on the steady-state locus, it turns out to be $K^{\star}$ since $G(K^{\star},\theta^{\star})=0$.
@@
$\mathrm{z}:=\operatorname{col}(K,\theta).$
@@
$\begin{bmatrix}\mathrm{w}\\ \chi\\ \mathrm{z}\end{bmatrix}\longmapsto\begin{bmatrix}\mathrm{w}\\ \tilde{\chi}\\ \tilde{\mathrm{z}}\end{bmatrix}:=\begin{bmatrix}\mathrm{w}\\ \begin{bmatrix}I_{{n}}&0\\ 0&\gamma I_{({n}+{m})(2{n}+{m})}\end{bmatrix}\left(\chi-\chi^{{\textnormal{ss}}}(\mathrm{w})\right)\\ \mathrm{z}-\begin{bmatrix}K^{\star}\\ \theta^{\star}\end{bmatrix}\end{bmatrix}.$ (29)
For notational convenience, we will sometimes refer to the components of $\tilde{\mathrm{z}}$ as $\operatorname{col}(\tilde{K},\tilde{\theta})$. Finally, the closed-loop dynamics (19) in the new coordinates (29) reads as
@@
where $\mathcal{A}(\tilde{\mathrm{z}}):=\mathcal{A}_{K}(\tilde{\mathrm{z}}_{1}+K^{\star})$ (cf. (25a)) and we introduced $h:\mathbb{R}^{n_{\mathrm{z}}}\times\mathbb{N}\to\mathbb{R}^{n_{\chi}}$, $g:\mathbb{R}^{n_{\chi}}\times\mathbb{R}^{n_{\mathrm{z}}}\times\mathbb{N}\to\mathbb{R}^{n_{\chi}}$, and $f:\mathbb{R}^{n_{\chi}}\times\mathbb{R}^{n_{\mathrm{z}}}\times\mathbb{N}\to\mathbb{R}^{n_{\mathrm{z}}}$ defined respectively as
$h(\tilde{\mathrm{z}},t)$ $:=\begin{bmatrix}{B}_{\star}\tilde{\mathrm{z}}_{1}\Pi_{\mathrm{x}}\mathrm{w}_{t}\\ 0_{({n}+{m})(2{n}+{m})}\end{bmatrix}$ (31a)
$g(\tilde{\chi},\tilde{\mathrm{z}},t)$ $:=\begin{bmatrix}0_{{n}}\\ \phi_{2}(\tilde{\chi}+\chi^{{\textnormal{ss}}}(\mathrm{w}_{t}),\tilde{\mathrm{z}}_{1}+K^{\star},\mathrm{w}_{t})-\phi_{2}(\chi^{{\textnormal{ss}}}(\mathrm{w}_{t}),K^{\star},\mathrm{w}_{t})\end{bmatrix}$ (31b)
$f(\tilde{\chi},\tilde{\mathrm{z}},t)$ $:=\begin{bmatrix}f_{1}(\tilde{\chi},\tilde{\mathrm{z}},t)\\ f_{2}(\tilde{\chi},\tilde{\mathrm{z}},t)\end{bmatrix},$ (31c)
with
$f_{1}(\tilde{\chi},\tilde{\mathrm{z}},t)$ $:=-G(\tilde{\mathrm{z}}_{1}+K^{\star},\tilde{\mathrm{z}}_{2}+\theta^{\star})$ (31d)
$f_{2}(\tilde{\chi},\tilde{\mathrm{z}},t)$ $:=\!-(\mathtt{unvec}\left(\tilde{\chi}_{2}\right)\!+\!H^{{\textnormal{ss}}}_{t})^{\dagger}\Big{(}(\mathtt{unvec}\left(\tilde{\chi}_{2}\right)\!+\!H^{{\textnormal{ss}}}_{t})\tilde{\mathrm{z}}_{2}+\mathtt{unvec}\left(\tilde{\chi}_{2}-\tilde{\chi}_{3}\right)\theta^{\star}\Big{)},$ (31e)
where, for the sake of readability, in (31) we used the shorthands $\tilde{\chi}:=\operatorname{col}(\tilde{\chi}_{1},\tilde{\chi}_{2},\tilde{\chi}_{3})$ and $\tilde{\mathrm{z}}=\operatorname{col}(\tilde{\mathrm{z}}_{1},\tilde{\mathrm{z}}_{2})$, we partitioned $\phi(\chi,K,\mathrm{w}):=\operatorname{col}(\phi_{1}(\chi,K,\mathrm{w}),\phi_{2}(\chi,K,\mathrm{w}))$ (cf. (25b)), we defined $H^{{\textnormal{ss}}}_{t}\in\mathbb{R}^{(n+m)\times(n+m)}$ as
@@
which represents the steady-state value of the state $H_{t}$ (see (20a) for the definition of $\mathrm{v}_{H}$) and we introduced the error coordinates $\tilde{H}\in\mathbb{R}^{(n+m)\times(n+m)}$ and $\tilde{S}\in\mathbb{R}^{(n+m)\times n}$, as
$\begin{bmatrix}\mathrm{w}\\ H^{\text{vc}}\\ S^{\text{vc}}\end{bmatrix}\longmapsto\begin{bmatrix}\mathrm{w}\\ \tilde{H}\\ \tilde{S}\end{bmatrix}:=\begin{bmatrix}\mathrm{w}\\ \mathtt{unvec}\left(H^{\text{vc}}-\mathrm{v}_{H}(\Pi_{H},\mathrm{w})\right)\\ \mathtt{unvec}\left(S^{\text{vc}}-\Pi_{S}\mathtt{vec}\left(\mathrm{w}\mathrm{w}^{\top}\right)\right)\end{bmatrix}.$ (33)
Some remarks are in order. We point out that with this transformation we obtained a dynamical system with two-time scales as the one described in Section 2.1 (cf. system (1)). As customary in the context of timescale separation theory, we distinguish between (i) the fast dynamics (30a) with state $\tilde{\chi}$, and (ii) the slow one (30b) with state $\tilde{\mathrm{z}}$. Figure 3 shows the mentioned interconnected structure of system (30).
Figure 3: Block diagram describing system (30).
It is also worth noting that, in this reformulation, the effect of the exogenous/dithering signal $\mathrm{w}_{t}$ has been embedded in the time dependency of $h$, $g$, and $f$. Finally, by using the definitions of $h$, $g$, and $f$ (cf. (31)) and the fact that $G(K^{\star},\theta^{\star})=0$ since $K^{\star}$ is the solution to problem (8), we note that
@@
for all $t\in\mathbb{N}$.